A non-toxic floor wax can be made from lemon juice and food-grade linseed oil. The amount of oil should be twice the amount of lemon juice. How much of each ingredient is needed to make 27 ounces of floor wax?
Let x represent the # of ounces of lemon juice and y represent the number of ounces of linseed oil.
Help!

Answer by checkley79(3341)

x+y=27
y=2x
x+2x=27
3x=27
x=27/3
x=9 oz. of lemon juice is needed.
y=2*9=18 oz. of oil needed.